An engineer wants to determine how the weight of a gas-powered car, x, affects gas mileage, y. The accompanying data represenI need assistance with the (a)

Please show details
Answer #1
weight(x) Miles per Gallon(y)   X - Mx Y - My    (X - Mx)2 (X - Mx)(Y - My)
3738 17 402.5455 -2.6364 162042.8 -1061.26
3770 17 434.5455 -2.6364 188829.8 -1145.62
2671 25 -664.455 5.3636 441499.8 -3563.89
3468 20 132.5455 0.3636 17568.3 48.1983
3325 20 -10.4545 0.3636 109.2975 -3.8017
2877 23 -458.455 3.3636 210180.6 -1542.07
3689 18 353.5455 -1.6364 124994.4 -578.529
2508 23 -827.455 3.3636 684681 -2783.26
3541 18 205.5455 -1.6364 42248.93 -336.347
3748 18 412.5455 -1.6364 170193.8 -675.074
3355 17 19.5455 -2.6364 382.0248 -51.5289

Sum of X = 36690
Sum of Y = 216
Mean X (MX) = 3335.4545
Mean Y (My) = 19.6364
Sum of squares (SSX) = 2042730.7273
Sum of products (SP) = -11693.1818

Regression Equation = ŷ = bX + a

b = SP/SSX = -11693.18/2042730.73 = -0.00572

a = MY - bMX = 19.64 - (-0.01*3335.45) = 38.72947

ŷ = -0.00572X + 38.73